# ASTRODEEP-JWST: NIRCam-HST multiband photometry and redshifts for half a million sources in six extragalactic deep fields

### v1

@@ Abstract
Aims. We present a set of photometric catalogs primarily aimed at providing the community with a comprehensive database for the study of galaxy populations in the high redshift Universe. The set gathers data from eight JWST NIRCam observational programs, targeting the Abell 2744 (GLASS-JWST, UNCOVER, DDT2756 and GO3990), EGS (CEERS), COSMOS and UDS (PRIMER), and GOODS North and South (JADES and NGDEEP) deep fields, for a total area of $\simeq$0.2 sq. degrees.
Methods. Photometric estimates are obtained by means of well-established techniques, including tailored improvements designed to enhance the performance on the specific dataset. We also include new measurements from HST archival data, thus collecting 16 bands spanning from 0.44 to 4.44 $\mu$m.
Results. A grand total of $\sim$530 thousand sources is detected on stacks of NIRCam 3.56 and 4.44 $\mu$m mosaics. We assess the photometric accuracy by comparing fluxes and colors against archival catalogs. We also provide photometric redshift estimates, statistically validated against a large set of robust spectroscopic data.
Conclusions. The catalogs are publicly available on the Astrodeep website.
@@ 1 Introduction
Two years since its first light, the James Webb Space Telescope (JWST, Gardner et al., 2006, 2023) has provided many cutting-edge quality data to many fields of astrophysical research. In particular, the study of the high redshift Universe has benefited from the joint effort of researchers designing tailored observational programs and exploiting their outcomes, with an outburst of activity within the very first weeks after the arrival of the first data (see e.g. Adamo et al., 2024, for a review).
@@
The unmatched depth and resolution of JWST infrared imaging and spectroscopy has enabled a wealth of analysis at intermediate and high redshifts that were beyond reach of previous instruments, from the measurement of optical rest-frame morphologies (e.g., Jacobs et al., 2023; Treu et al., 2023; Kartaltepe et al., 2023), stellar masses (e.g., Santini et al., 2023; Weibel et al., 2024; Wang et al., 2024) and star-formation histories (e.g., Dressler et al., 2023; Looser et al., 2023; Ciesla et al., 2023) of galaxies up to $z\simeq 10$, to the abundance and properties of red and optically-dark sources (e.g., Glazebrook et al., 2023; Kirkpatrick et al., 2023; Pérez-González et al., 2023a; Rodighiero et al., 2024) and quiescent galaxies (e.g., Carnall et al., 2023; Nanayakkara et al., 2024; Wright et al., 2024; Ward et al., 2024) up to $z\sim 6$, the properties of compact red sources (Pérez-González et al., 2024; Williams et al., 2024; Kokorev et al., 2024), or the identification of extreme emission line galaxies in the reionization epoch (e.g. Davis et al., 2023). In particular, JWST has challenged our view of the early epochs of the cosmos. A number of exciting discoveries on the first phases of galaxy formation and evolution have resulted in more questions than answers, as we have yet to frame and understand our observations within a fully consistent theoretical framework. The most consolidated result so far is the evidence of a striking overabundance of bright galaxies at $z\gtrsim 9-10$ compared to most predictions (e.g., Castellano et al., 2022, 2023; Naidu et al., 2022; Finkelstein et al., 2023a, b; McLeod et al., 2024; Pérez-González et al., 2023b; Carniani et al., 2024). These distant sources have been mostly identified by means of color selections or spectral energy distribution (SED) fitting, and the result is still a matter of debate, with many possible explanations proposed (e.g. Ferrara et al., 2023; Dekel et al., 2023; Mason et al., 2023; Trinca et al., 2024; Padmanabhan & Loeb, 2023; Yung et al., 2024). While spectroscopic follow-up is crucial to understand the physical processes at play, and indeed they currently seem to confirm the early results (e.g. Arrabal Haro et al., 2023a; Harikane et al., 2024; Castellano et al., 2024), photometric data still constitute the primary way to collect statistically significant samples, beyond providing targets for spectroscopy.
@@
In this paper we present our analysis of eight deep-sky NIRCam observational programs: CEERS, DDT2756, GLASS-JWST, GO3990, JADES, NGDEEP, PRIMER and UNCOVER, from which we created a new set of photometric catalogs, mainly finalized to provide a consistent database for the study of the early phases of galaxy evolution in the high redshift Universe ($z\geq 3$). The programs target six among the most well-known and studied areas of the sky, already observed with HST and ground-based facilities (in particular, the CANDELS and Frontier Fields campaigns, see Grogin et al., 2011; Koekemoer et al., 2011; Lotz et al., 2014), and subject of many ground-breaking studies in the past decades: (i) the GLASS-JWST, UNCOVER, DDT2756 and GO3990 programs cover the Abell 2744 cluster of galaxies, also known as Pandora’s cluster, and its surrounding area. For simplicity, we will refer to this extended region as ABELL2744; (ii) the CEERS survey overlaps with the Extended Groth Strip (EGS) field (Davis et al., 2007); (iii-iv) the PRIMER program covers two areas, one overlapping with the UKIDSS Ultra-deep Survey field (UDS, Lawrence et al., 2007) and the other with the COSMOS field (Nayyeri et al., 2017); (vi) JADES and NGDEEP overlap with the GOODS-North and ECDFS/GOODS-South regions (Giacconi et al., 2002; Giavalisco et al., 2004). The total area observed by these programs is $\simeq$0.2 sq. degrees. While some catalogs for these observations are already available (e.g. Paris et al., 2023; Rieke et al., 2023; Weaver et al., 2024), the aim of this new dataset is to provide a large, self-consistent database mainly aimed at the study of the high redshift Universe. Therefore, we mostly focused our attention on the detection of faint sources when choosing detection parameters. Unsurprisingly, this choice can lead to some degree of tension when comparing our results to other catalogs (see Sect. 4).
@@
The new catalogs are obtained with well-tested algorithms and techniques, largely building upon previous releases of our group (Merlin et al., 2022; Paris et al., 2023, respectively M22 and P23 hereafter), but with substantial improvements, that we describe in Section 3. Up-to-date data and calibration files have been used whenever available (detailed information is provided). We also estimated photometric redshift, which we obtained using the zphot SED-fitting software, first described in Fontana et al. (2000), and EAzY (Brammer et al., 2008) exploiting three different sets of templates. Predictably, we find that while for high signal-to-noise (SNR) the four runs typically yield estimates in reasonable agreement, faint objects are less well constrained, often resulting in divergent fits. Finally, we validate our results comparing them to the literature. We find substantial agreement with other photometric catalogs, and good statistics checking the photo-$z$’s against spectroscopic data.
@@
The paper is organized as follows. In Sect. 2 we describe the dataset. In Sect. 3 we summarise the adopted techniques, referencing previous publications when useful, and highlighting the differences with previous works. In Sect. 4 we discuss the validation of our new catalogs against published data, considering direct comparisons of colors. In Sect. 5 we discuss photometric redshifts. In Sect. 6 we make some final remarks and conclusions. The catalog format is described in Appendix A. We use AB magnitudes (Oke & Gunn, 1983) and we assume a $\Lambda$CDM cosmology ($H_{0}=70$ km/s/Mpc, $\Omega_{m}=0.27$ and $\Omega_{\Lambda}=0.73$). The catalogs are publicly available on the Astrodeep website (https://astrodeep.eu).
@@ 2 Dataset
In this Section we summarize the properties of the dataset we used to obtain the set of photometric catalogs.
For all the fields but ABELL2744, we collected the NIRCam 30mas mosaics created by the teams of each program. In most cases, ancillary HST images were also made available; when needed, we re-projected them on the same grid of the NIRCam mosaics, and checked for astrometric consistency (see Sect. 6). All of the images were also scaled to $\mu$Jy units, so that AB magnitudes can be obtained from the flux measurements applying a constant zero-point (ZP) of 23.9. In the following we detail the process case by case; because of the composite nature of the dataset, which comprises images reduced by different teams at different times, the specific parameters of the reduction processes were unavoidably non-uniform. The different calibration files and pipeline versions that have been used are summarized in Table 1.
In order to create a formally homogeneous set of catalogs, we selected a fixed set of pass-band filters, thus also facilitating the photo-$z$ estimates for which tailored libraries of models are required. The chosen bands are: HST ACS F435W, F606W, F775W and F814W; HST WFC3 F105W, F125W, F140W and F160W; and NIRCam F090W, F115W, F150W, F200W, F277W, F356W, F410M and F [^4] W. The main features of these pass-bands are summarized in Table 2. Note that not all of the bands are available in all fields; we provide the relevant details in the following subsections. The areas given for each field are indicative, as many pass-bands have different sky coverage; the values refer to the F [^4] W band area.
@@
PRIMER-UDS 0.6 11.17.0 1118*** 1.10.2
*GLASS-JWST, UNCOVER, DDT2756 and GO3390
**CEERS full field mosaics are constructed from two
separate public data releases: DR0.5 for NIRCam
pointings 1,2,3,6 and DR0.6 for NIRCam pointings
4,5,7,8,9,10 (DOI 10.17909/z7p0-8481)
***pmap 1117 for F356W, F410M, F [^4] W
Table 2: Main features of the pass-band filters included in the catalogs. $fr$ is the fraction of the flux within a circular aperture of diameter 0.2”. Wavelengths are given in nanometers; FWHMs are given in arcseconds.
@@ 2.1 ABELL2744
Imaging data from four programs have been combined in a single set of mosaics: GLASS-JWST (ERS 1324, P.I. Treu; no F410M, Treu et al., 2022), UNCOVER (GO 2561, P.I. Labbé; no F090W, Bezanson et al., 2022), DDT 2756 (P.I. Chen; no F090W and F410M), and GO 3990 (P.I. Morishita; no F410M). The resulting combined field of view (FoV) is centered on the galaxy cluster, and covers an area of $\simeq$45.7 sq. arcmin. With respect to P23, new data have been received and added: the GO 3990 images in the UNCOVER region, and a new set of observations of the GLASS-JWST region acquired in July 2023 to correct the original 2022 images which were affected by a wing-tilt event in the short-wavelength bands (SW hereafter, i.e. F090W, F115W, F150W and F200W; the redder bands are denoted long-wavelength, LW hereafter). The reduction has been re-done from scratch with new calibration files, following the procedure described in P23. The raw uncal images have been retrieved from the MAST archive [^1] https://mast.stsci.edu/, and combined to cal images by applying the first two stages of the official $JWST$ calibration pipeline (calwebb_detector1 and calwebb_image2) with the latest calibration and reference files available to date, see Tab. 1. We then applied our modified version of the official pipeline using a number of custom procedures developed by our team to correct for defects – snowballs, non-linear pixels, 1/$f$-noise, wisps and claws – and finally combined the calibrated images into mosaics and ancillary weight and error (root mean square, RMS) maps (see P23 Sect. 2.2.1 for details). We then complemented the NIRCam dataset with archival HST mosaics, reduced and publicly released by G. Brammer [^2] https://s3.amazonaws.com/grizli-v2/JwstMosaics/
@@
We point out that this field is different from all the others, in that it is centered on a galaxy cluster. Working on this same region for the Frontier Fields campaign, Merlin et al. (2016b) developed a sophisticated technique to accurately subtract bright foreground galaxies and intra-cluster light from all the analysed bands. However, such a technique is complex and time-consuming, so we postpone its application to future work. As mentioned in P23, the global background and 1/$f$-noise subtraction techniques effectively removes most of the intra-cluster light from the images. Furthermore, we provide photometric estimates including local background subtraction (see Sect. 3.4), which effectively removes residual intra-cluster light while also mitigating spurious effects created by the global processing. Nevertheless, we warn the users that the sources close to the center of the cluster should be treated with caution (see also Sect. 5); also, potential very high redshift objects magnified by (but close to) the cluster center are currently impossible to detect.
@@ 2.2 CEERS
Data from the CEERS program (ERS 1345, P.I. Finkelstein; no F090W, Finkelstein et al., 2022) cover an area of $\simeq$94.6 sq. arcmin. Images have been reduced by M. Bagley combining two epochs of observations (see Bagley et al., 2023). The individual pointings are available from CEERS public releases 0.5 and 0.6333https://ceers.github.io/dr05.html, https://ceers.github.io/dr06.html, and at MAST as High Level Science Products via DOI 10.17909/z7p0-8481. We have drizzled all ten individual pointings from these two releases into single mosaics for this paper. For HST we used the EGS dataset from CANDELS (Stefanon et al., 2017, no F435W and F775W data), with the addition of the F105W band from the CEERS HDR1 reduction [^4] https://ceers.github.io/releases#hdr1. A new catalog on a more recent, improved reduction of the NIRCam data, and obtained applying slightly different techniques, is going to be published soon by the CEERS team (Cox et al., in preparation).
@@ 2.3 JADES and NGDEEP
Data from the JADES program (GTO 1180, P.I. Eisenstein, and GTO 1210, P.I. Luetzgendorf, Eisenstein et al., 2023) cover an area of $\simeq$83.0 sq. arcmin in the GOODS-North region (JADES-GN hereafter), and of $\simeq$84.5 sq. arcmin in the GOODS-South region (JADES-GS hereafter). The images are available to the public and we used the v2.0 version for JADES-GS and the v1.0 version for JADES-GN. We complemented the NIRCam data using the Hubble Legacy Fields images (Illingworth et al., 2016).
The NGDEEP program (P.I. Finkelstein; no F090W and F410M, Bagley et al., 2024) adds an area of $\simeq$9.5 sq. arcmin from the outer ECDFS area (with a marginal overlap with the GOODS-South field). Since the observed FoV does not overlap with JADES-GS, we kept the two fields separated, creating two catalogs and analysing them individually. We used the same imaging data of Leung et al. (2023), which only includes the first epoch of observation. This first epoch suffered from a lack in depth due to the DEEP8 readout pattern with a small number of groups. An updated catalog based on both significant improvements on the first epoch and including the second epoch will be published soon (Leung et al., in preparation).

### v3

# ASTRODEEP-JWST: NIRCam-HST multi-band photometry and redshifts for half a million sources in six extragalactic deep fields

@@ Abstract
Aims. We present a set of photometric catalogues primarily aimed at providing the community with a comprehensive database for the study of galaxy populations in the high-redshift Universe. The set gathers data from eight JWST NIRCam observational programs, targeting the Abell 2744 (GLASS-JWST, UNCOVER, DDT2756, and GO3990), EGS (CEERS), COSMOS and UDS (PRIMER), and the GOODS North and South (JADES and NGDEEP) deep fields. This dataset covers a total area of $\simeq$0.2 sq. degrees.
Methods. We obtained photometric estimates by means of well-established techniques, including tailored improvements designed to enhance the performance on the specific dataset. We also included new measurements from HST archival data, spanning 16 bands from 0.44 to 4.44 $\mu$m.
Results. A grand total of $\sim$530 thousand sources were detected on stacks of NIRCam 3.56 and 4.44 $\mu$m mosaics. We assessed the photometric accuracy by comparing fluxes and colours against archival catalogues. We also provide photometric redshift estimates, statistically validated against a large set of robust spectroscopic data.
Conclusions. The catalogues are publicly available on the Astrodeep website.
@@ 1 Introduction
Two years since its first light, the James Webb Space Telescope (JWST; Gardner et al., 2006, 2023) has provided a vast store of cutting-edge, quality data to many fields of astrophysical research. In particular, the study of the high-redshift Universe has benefitted from the joint effort of researchers designing tailored observational programmes and exploiting their outcomes, with an outburst of activity within the very first weeks after the arrival of the first data (see e.g. Adamo et al., 2024, for a review).
@@
The unmatched depth and resolution of JWST infrared imaging and spectroscopy has enabled a wealth of analysis at intermediate and high redshifts that were beyond the reach of previous instruments. Examples include the measurement of optical rest-frame morphologies (e.g. Ferreira et al., 2022, 2023; Jacobs et al., 2023; Treu et al., 2023; Kartaltepe et al., 2023), stellar masses (e.g. Santini et al., 2023; Weibel et al., 2024; Wang et al., 2024a), and star-formation histories (e.g. Dressler et al., 2023; Looser et al., 2023; Ciesla et al., 2024; Conselice et al., 2024) of galaxies up to $z\simeq 10$, to the abundance and properties of red and optically dark sources (e.g. Glazebrook et al., 2023; Kirkpatrick et al., 2023; Pérez-González et al., 2023a; Rodighiero et al., 2024) as well as quiescent galaxies (e.g. Carnall et al., 2023; Nanayakkara et al., 2024; Wright et al., 2024; Ward et al., 2024) up to $z\sim 6$, the properties of compact red sources (Pérez-González et al., 2024; Williams et al., 2024; Kokorev et al., 2024), and the identification of extreme emission line galaxies in the epoch of reionization (e.g. Davis et al., 2023). In particular, JWST has challenged our view of the early epochs of the cosmos. A number of exciting discoveries on the first phases of galaxy formation and evolution have resulted in more questions than answers, as we have yet to frame and understand our observations within a fully consistent theoretical framework. The most consolidated result so far is the evidence of a striking overabundance of bright galaxies at $z\gtrsim 9-10$ compared to most predictions (e.g. Castellano et al., 2022, 2023; Naidu et al., 2022; Finkelstein et al., 2023, 2024; McLeod et al., 2024; Pérez-González et al., 2023b; Carniani et al., 2024). These distant sources have been mostly identified by means of colour selections or spectral energy distribution (SED) fitting, and the result is still a matter of debate, with numerous possible explanations proposed (e.g. Ferrara et al., 2023; Dekel et al., 2023; Mason et al., 2023; Trinca et al., 2024; Padmanabhan & Loeb, 2023; Yung et al., 2024; Harvey et al., 2024). While spectroscopic follow-up is crucial to understand the physical processes at play – and, indeed, they currently seem to confirm the early results (e.g. Arrabal Haro et al., 2023a; Harikane et al., 2024; Castellano et al., 2024), photometric data still constitute the primary way to collect statistically significant samples, beyond providing targets for spectroscopy.
@@
In this paper, we present our analysis of eight deep-sky NIRCam observational programmes: CEERS, DDT2756, GLASS-JWST, GO3990, JADES, NGDEEP, PRIMER, and UNCOVER. We used these data to create a new set of photometric catalogues, mainly finalised to provide a consistent database for the study of the early phases of galaxy evolution in the high-redshift Universe ($z\geq 3$). The programmes target six of the most well-known and studied areas of the sky, which have already been observed with HST and ground-based facilities (in particular, the CANDELS and Frontier Fields campaigns; see Grogin et al., 2011; Koekemoer et al., 2011; Lotz et al., 2017), and have been subject of many ground-breaking studies in the past decades: (i) the GLASS-JWST, UNCOVER, DDT2756 and GO3990 programmes cover the Abell 2744 cluster of galaxies, also known as Pandora’s cluster, and its surrounding area. For simplicity, we will refer to this extended region as ABELL2744; (ii) the CEERS survey overlaps with the Extended Groth Strip (EGS) field (Davis et al., 2007); (iii-iv) the PRIMER programme covers two areas, one overlapping with the UKIDSS Ultra-deep Survey field (UDS, Lawrence et al., 2007) and the other with the COSMOS field (Nayyeri et al., 2017); (vi) JADES and NGDEEP overlap with the GOODS-North and ECDFS/GOODS-South regions (Giacconi et al., 2002; Giavalisco et al., 2004). The total area observed by these programmes is $\simeq$0.2 sq. degrees. While some catalogues for these observations are already available (e.g. Paris et al., 2023; Rieke et al., 2023; Weaver et al., 2024), the aim of this new dataset is to provide a large, self-consistent database mainly aimed at the study of the high-redshift Universe. Therefore, we mostly focused our attention on the detection of faint sources when choosing detection parameters. Unsurprisingly, this choice can lead to some degree of tension when comparing our results to other catalogues (see Sect. 4).
@@
The new catalogues are obtained with well-tested algorithms and techniques, largely building upon previous releases of our group (Merlin et al., 2022; Paris et al., 2023, respectively M22 and P23 hereafter), but with substantial improvements, which we describe in Sect. 3. Up-to-date data and calibration files have been used whenever available (detailed information is provided). We also estimated photometric redshifts, which we obtained using the SED-fitting software packages zphot, first described in Fontana et al. (2000), and EAzY (Brammer et al., 2008), for which we used three different sets of templates. Predictably, comparing the results from these four runs we find that they typically are in good agreement for high signal-to-noise (S/N) sources, while faint objects are less well constrained, often resulting in divergent fits. Finally, we validated our results comparing them to the literature. We find substantial agreement with other photometric catalogues and good statistics when checking the photo-$z$ results against the spectroscopic data.
@@
The paper is organised as follows. In Sect. 2 we describe the dataset. In Sect. 3 we summarise the adopted techniques, referencing previous publications when useful and highlighting the differences with previous works. In Sect. 4 we discuss the validation of our new catalogues against published data, considering direct comparisons of colours. In Sect. 5, we discuss the photometric redshifts. In Sect. 6, we make some final remarks and conclusions. The catalogue format is described in Appendix A (available on Zenodo; see Sect. 7). We use AB magnitudes (Oke & Gunn, 1983) and we assume an $\Lambda$CDM cosmology ($H_{0}=70$ km/s/Mpc, $\Omega_{m}=0.27,$ and $\Omega_{\Lambda}=0.73$).
@@ 2 Dataset
In this section. we summarise the properties of the composite dataset we used to obtain the photometric catalogues presented in this work. For all the fields but ABELL2744, we collected the NIRCam 30mas mosaics created by the teams of each programme. In most cases, ancillary HST images were also made available; when needed, we re-projected them on the same grid of the NIRCam mosaics, and checked for astrometric consistency (see Sect. 3.1). All of the images were also scaled to $\mu$Jy units, so that AB magnitudes can be obtained from the flux measurements applying a constant zero-point (ZP) of 23.9. In the following, we detail the process case by case; because of the composite nature of the dataset, which comprises images reduced by different teams at different times, the specific parameters of the reduction processes were unavoidably non-uniform. The different calibration files and pipeline versions used in this study are summarised in Table $1$See the official JWST pipeline web-page for detailed information on the keywords, https://jwst-pipeline.readthedocs.io/en/latest/.
In order to create a formally homogeneous set of catalogues, we selected a fixed set of pass-band filters, thereby also facilitating the photo-$z$ estimates for which tailored libraries of models are required. The chosen bands are: HST ACS F435W, F606W, F775W and F814W; HST WFC3 F105W, F125W, F140W and F160W; and NIRCam F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F [^4] W. The main features of these pass-bands are summarised in Table 2. We point out that not all of the bands are available in all fields; we provide the relevant details in the following subsections. The areas given for each field are indicative, as many pass-bands have different sky coverage; the values refer to the F [^4] W band area.
@@
[…]
222
Table 2: Main features of the pass-band filters included in the catalogues.
@@ 2.1 ABELL2744
Imaging data from four programmes were combined in a single set of mosaics: GLASS-JWST (ERS 1324, P.I. Treu; no F410M, Treu et al., 2022), UNCOVER (GO 2561, P.I. Labbé; no F090W, Bezanson et al., 2022), DDT 2756 (P.I. Chen; no F090W and F410M), and GO 3990 (P.I. Morishita; no F410M). The resulting combined field of view (FoV) is centred on the galaxy cluster, and covers an area of $\simeq$45.7 sq. arcmin. With respect to P23, new data have been received and added: the GO 3990 images in the UNCOVER region and a new set of observations of the GLASS-JWST region acquired in July 2023 to correct the original 2022 images that were affected by a wing-tilt event in the short-wavelength bands (SW hereafter, i.e. F090W, F115W, F150W, and F200W; the redder bands are denoted long-wavelength, LW hereafter). The reduction has been re-done from scratch with new calibration files, following the procedure described in P23. The raw uncal images have been retrieved from the MAST archive [^4] https://mast.stsci.edu/, and combined to cal images by applying the first two stages of the official JWST calibration pipeline (calwebb_detector1 and calwebb_image2) with the latest calibration and reference files available to date (see Table 1). We then applied our modified version of the official pipeline using a number of custom procedures developed by our team to correct for defects: ‘snowballs’, non-linear pixels, 1/$f$-noise, ‘wisps’ and ‘claws’. We aligned the calibrated images to Gaia-DR3 astrometry, and finally we combined them into mosaics and ancillary weight and error (root mean square, RMS) maps, (see P23 Sect. 2.2.1 for details). We then complemented the NIRCam dataset with archival HST mosaics, reduced, and publicly released by G. Brammer [^5] https://s3.amazonaws.com/grizli-v2/JwstMosaics/
@@
. A global background subtraction on all the mosaics was performed with SExtractor (Bertin & Arnouts, 1996).
We point out that this field is different from all the others because it is centred on a galaxy cluster. Working on this same region for the Frontier Fields campaign, Merlin et al. (2016b) developed a sophisticated technique to accurately subtract bright foreground galaxies and intra-cluster light from all the analysed bands. However, such a technique is complex and time-consuming, so we postpone its application to future work. As mentioned in P23, the global background and 1/$f$-noise subtraction techniques effectively remove most of the intra-cluster light from the images. Furthermore, we provide photometric estimates including local background subtraction (see Sect. 3.4), which effectively removes residual intra-cluster light while also mitigating spurious effects created by the global processing. Nevertheless, we warn users that the sources close to the centre of the cluster should be treated with caution, since their photometry might be affected by the residual light of the cluster members or by artefacts caused by the global background subtraction (see also Sect. 3.4.2). Also, potential very-high-redshift objects magnified by (but close to) the cluster centre are currently impossible to detect.
@@ 2.2 CEERS
Data from the CEERS programme (ERS 1345, P.I. Finkelstein; no F090W, Finkelstein et al., 2022) cover an area of $\simeq$94.6 sq. arcmin. Images have been reduced by M. Bagley combining two epochs of observations (see Bagley et al., 2023). The individual pointings are available from CEERS public releases 0.5 and 0.6666https://ceers.github.io/dr05.html, https://ceers.github.io/dr06.html and at MAST as High Level Science Products via DOI 10.17909/z7p0-8481. We have drizzled all ten individual pointings from these two releases into single mosaics for this paper. For HST we used the EGS dataset from CANDELS (Stefanon et al., 2017, no F435W and F775W data), with the addition of the F105W band from the CEERS HDR1 reduction [^7] https://ceers.github.io/releases#hdr1. A new catalogue based on a more recent, improved reduction of the NIRCam data, and obtained applying slightly different techniques, is going to be published soon by the CEERS team (Cox et al., in preparation).
@@ 2.3 JADES and NGDEEP
Mosaics from the JADES programme (GTO 1180, P.I. Eisenstein, and GTO 1210, P.I. Luetzgendorf, Eisenstein et al., 2023), including additional data from the FRESCO programme (PID 1895, P.I. Oesch), cover an area of $\simeq$83.0 sq. arcmin in the GOODS-North region (JADES-GN hereafter), and of $\simeq$84.5 sq. arcmin in the GOODS-South region (JADES-GS hereafter). The images are available to the public. We used the v2.0 version for JADES-GS and the v1.0 version for JADES-GN. We complemented the NIRCam data using the Hubble Legacy Fields images (Illingworth et al., 2016).
The NGDEEP programme (P.I. Finkelstein; no F090W and F410M, Bagley et al., 2024) adds an area of $\simeq$9.5 sq. arcmin from the outer ECDFS area, with a marginal overlap with the GOODS-South field. Since the observed FoV does not overlap with JADES-GS, we kept the two fields separated, creating two catalogues and analysing them individually. We used the same imaging data of Leung et al. (2023), which only includes the first epoch of observation. This first epoch suffered from a lack of depth due to the DEEP8 readout pattern with a small number of groups. An updated catalogue based on both significant improvements to the first epoch and including the second epoch will be published in the near future (Leung et al., in preparation).